OpenStudy (anonymous):

differentiate f(x)=e^x/9x^2

OpenStudy (anonymous):

use the improved version of the quotient rule: (u/(v^n))'=(vu'-nuv')/(v^n+1)

OpenStudy (anonymous):

f'(x) =[e^x (9x^2) -18x(e^x)]/(81x^4)

OpenStudy (anonymous):

should be .../(v^(n+1))

OpenStudy (anonymous):

we may simlify then.

OpenStudy (anonymous):

= e^x(x-2)/9x^3

OpenStudy (anonymous):

improved version of the quotient rule: (u/(v^n))'=(vu'-nuv')/(v^n+1) is the simplification built in in the formula